OpenStudy (anonymous):

Solve this system and list the method that you chose (graphing, substitution or elimination). y = 2x – 12 y = -x + 3

OpenStudy (anonymous):

Easiest way to me is to put them together and solve. So: -x + 3 = 2x - 12 +x +x 3 = 3x - 12 +12 +12 15 = 3x Continue to solve for x